If we have 200.0mL of gas at 17.0 C what will be the volume of this gas at 10.0 C


1
Expert's answer
2021-01-11T03:53:51-0500

V1 = 200.0 mL = 0.2 L

T1 = 17.0oC + 273 = 290 K

V2 = unknown

T2 = 10.0oC + 273 = 283 K

P1 = P2 = const


Solution:

Since the pressure doesn't change, Charles' Law can be used.

Charles' Law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the final volume, solve the equation for V2:

V2 = V1T2/T1

V2 = (0.2 L × 283 K) / (290 K) = 0.1952 L = 195.2 mL

V2 = 195.2 mL


Answer: V2 = 195.2 mL